OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (mathmale):

Joe: Please look up the slope formula if you don't already have it in front of you. Find the slope the line connecting these two points. Then, choose either of the two points. Substitute its x- and y-coordinates, along with your slope, m, into the point-slope formula for the equation of a straight line: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

Finally, re-write your result to resemble this form: Ax + By = C (standard form).

OpenStudy (anonymous):

I still don't get it. so lets try this together so 9-3=6